OpenStudy (anonymous):

Find The Sum 4 2 - + - 7 5

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}=\frac{a\times d+b\times c}{b\times d}\]is how you add

OpenStudy (anonymous):

you get \[\frac{4\times 5+7\times 2}{7\times 5}\]

OpenStudy (anonymous):

Yes Thanks

OpenStudy (anonymous):

yw